Is this as simple as adding two numbers togather? Find 2 consecutive even integers that have a sum of 450.
Answer by jojo14344(1513)

First, let put "x' as the 1st integer.
Since they are consecutive even integer, the # should follow this should be "x+2" right?(2nd integer)
Therefore, 1st int.+2nd int.=450
x+x+2=450
2x=448
x=224, the 1st integer
And, "x+2"=224+2=226, the 2nd integer.
The sum of the two, 224+226=450, cool!
Thank you,
Jojo